OpenStudy (anonymous):

Solve the systems using substitution. 6x + 2y = 0 4x + y = -1

OpenStudy (anonymous):

here's one substitution you can do: in the second equation, let y=-4x-1. substitute this into the first equation: 6x + 2(-4x-1) = 0 can you take it from here?

OpenStudy (anonymous):

No :/ Please show me step by step

OpenStudy (anonymous):

No :/ Please show me step by step

OpenStudy (anonymous):

6x + 2(-4x-1) = 0 carry out distributive property on the multiplication with the parenthesis: 6x + 2(-4x-1) = 0 6x + (2*-4x) - (2*-1) = 0 6x + (-8x) - (-2) = 0 how 'bout now?.... can you do the rest?

OpenStudy (anonymous):

6x + 2 (-4x + 1) = 0 6x – 8x + 2 = 0 -2 + 2 = 0

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Whoops I thought it was a positive 1 I didnt see the negative

OpenStudy (anonymous):

6x + (-8x) - (-2) = 0 6x - 8x + 2 = 0 -2x + 2 = 0 how 'bout from here then....

OpenStudy (anonymous):

X = -1

OpenStudy (anonymous):

is that right?